# Galaxy assembly bias in cosmological hydrodynamical simulations – a comparison between SIMBA and IllustrisTNG

### v1

@@ Abstract
pubyear: 2025pagerange: Galaxy assembly bias in cosmological hydrodynamical simulations – a comparison between SIMBA and IllustrisTNGGalaxy assembly bias in cosmological hydrodynamical simulations – a comparison between SIMBA and IllustrisTNG
@@ 1 Introduction
The spatial distribution of galaxies is a key observable for cosmological studies, as it traces the underlying large-scale structure (LSS) of the Universe, which is primarily composed of invisible dark matter. In the framework of hierarchical structure formation, galaxies form within dark matter haloes that originate from primordial density fluctuations and grow through gravitational collapse, accretion, and mergers (White & Rees, 1978; Blumenthal et al., 1984; Lacey & Cole, 1993). A precise understanding of the galaxy–halo connection is therefore essential for interpreting data from current and forthcoming galaxy redshift surveys, such as the Dark Energy Spectroscopic Instrument (DESI; e.g., DESI Collaboration et al., 2024, 2025a, 2025b, 2025c) and the Euclid mission (e.g., Euclid Collaboration et al., 2022, 2025).
@@
One of the earliest approaches to modelling the galaxy–halo connection is the halo occupation distribution (HOD; e.g., Benson et al., 2000; Seljak, 2000; Peacock & Smith, 2000; White et al., 2001; Berlind & Weinberg, 2002; Cooray & Sheth, 2002), which describes the average number of galaxies hosted by a dark matter halo as a function of its mass. The simplest form of the HOD assumes that halo mass alone determines galaxy occupancy, a notion motivated by excursion set theory (Press & Schechter, 1974; Bond et al., 1991), which predicts that halo properties beyond mass are uncorrelated with the cosmic environment due to its random-walk formalism (Lemson & Kauffmann, 1999). However, this assumption was later challenged by Sheth & Tormen (2004), who demonstrated using $N$-body simulations that haloes in dense environments tend to form earlier than haloes of the same mass in less dense regions. Building on this, Gao et al. (2005) used the Millennium Simulation (Springel et al., 2005) to show that halo clustering depends not only on mass but also on halo formation time. This secondary dependence was subsequently extended to other halo properties, such as concentration and spin (Wechsler et al., 2006; Gao & White, 2007; Angulo et al., 2008) and the cosmic web environment (Borzyszkowski et al., 2017; Tojeiro et al., 2017; Yang et al., 2017; Ramakrishnan et al., 2019). A dependence of halo clustering on properties beyond their mass is commonly referred to as halo assembly bias (HAB; e.g. Wechsler & Tinker, 2018), although the secondary properties may not correspond exactly to the halo assembly history.
@@
If galaxy formation is influenced by the secondary halo properties that give rise to HAB, then the core assumption of the simplest HOD model – that galaxy occupancy depends solely on halo mass – is violated. This violation can systematically alter predictions of galaxy clustering. The phenomenon in which the galaxy–halo connection (see Wechsler & Tinker, 2018 for a review) depends on halo properties beyond mass is known as galaxy assembly bias (GAB).
@@
The existence of GAB was confirmed shortly after the discovery of HAB. Zhu et al. (2006) showed that galaxy properties and occupancy correlate with secondary halo properties. Croton et al. (2007) introduced the shuffling technique for quantifying GAB, and used it to demonstrate that galaxy clustering is sensitive to halo properties beyond mass. Since then, numerous studies have examined the impact of GAB on galaxy clustering and the degree to which it is driven by specific halo internal properties such as age, concentration, or spin (Zu et al., 2008; Chaves-Montero et al., 2016; Zehavi et al., 2018; Artale et al., 2018; Contreras et al., 2019; Xu et al., 2021; García-Moreno & Chaves-Montero, 2025) – or by environmental properties such as smoothed overdensity and tidal anisotropy (Zehavi et al., 2018; Artale et al., 2018; Xu et al., 2021; Alam et al., 2023; Montero-Dorta & Rodriguez, 2024; Wang et al., 2024; Perez et al., 2024). These investigations have primarily relied on galaxy catalogues generated from cosmological hydrodynamical simulations (mainly IllustrisTNG; Nelson et al., 2019) and/or semi-analytical galaxy formation models applied to $N$-body simulations (e.g., Guo et al., 2013).
@@
Despite this extensive body of work, the physical origin of GAB remains unclear. Moreover, its presence in observational data is still debated (e.g., Lin et al., 2016; Miyatake et al., 2016; Vakili & Hahn, 2019; Salcedo et al., 2022; Contreras et al., 2023; Zhang et al., 2025). If GAB is indeed present in real data, then halo mass alone is insufficient to fully describe the galaxy–halo connection in the era of precision cosmology (Wu et al., 2008; Zentner et al., 2014; McCarthy et al., 2019; Hadzhiyska et al., 2020; Wang et al., 2025). Understanding the underlying physical mechanism of GAB is therefore both critical and timely for interpreting data from galaxy redshift surveys.
@@
In this paper, we propose that comparing the strength and evolution of GAB across different cosmological hydrodynamical simulations can provide valuable insight into its physical origin. Previous studies have shown that GAB signals in IllustrisTNG and in semi-analytical models are consistent (e.g., Hadzhiyska et al., 2021), but direct comparisons between different hydrodynamical simulations have not been made. Although modern hydrodynamical simulations broadly reproduce observed galaxy properties, they differ significantly in aspects such as gas heating efficiency and small-scale clustering (e.g., Ni et al., 2023). Comparing GAB across simulations with differing baryonic prescriptions can thus help isolate the physical processes responsible for the emergence of GAB.
@@ 2 Data
We use two state-of-the-art cosmological hydrodynamical simulations to study the redshift evolution of GAB: SIMBA (Davé et al., 2019) and IllustrisTNG (Nelson et al., 2019). These simulations include physical prescriptions for various baryonic processes, such as stellar winds, gas cooling, and AGN feedback. Importantly, the baryonic physics implemented in SIMBA and TNG differ in key aspects, such as the treatment of gas heating by feedback (see, e.g., Ni et al., 2023). Despite these differences, both simulations reproduce the observed galaxy stellar mass function at $z=0$ with high fidelity. Comparing the GAB evolution in the two simulations thus provides an opportunity to gain insights into the physical mechanisms driving GAB. Below, we briefly summarise the relevant features of each simulation.
@@ 2.1 SIMBA
The SIMBA simulation suite models galaxy formation within a cosmological context using a modified version of the GIZMO code (Hopkins, 2015, 2017) with the meshless finite mass (MFM) hydrodynamics solver. SIMBA incorporates a range of baryonic feedback processes, including stellar feedback, AGN feedback in the form of radiative winds and collimated jets, and X-ray feedback from black hole accretion (Davé et al., 2019).
@@ 2.2 IllustrisTNG
The IllustrisTNG project comprises a suite of large-volume cosmological magnetohydrodynamical simulations (Weinberger et al., 2017; Pillepich et al., 2018) based on the moving-mesh code AREPO (Springel, 2010). We use the largest run of the IllustrisTNG project, namely TNG300-1 (Nelson et al., 2019; Pillepich et al., 2017), which we refer to as TNG hereafter for simplicity. TNG simulates a cubic volume with $L_{\mathrm{box}}=205\,h^{-1}{\rm Mpc}$ using $2500^{3}$ particles for both dark matter and baryons. The mass resolution for baryon and dark matter are $7.6\times 10^{6}\,h^{-1}\,\mathrm{M_{\odot}}$ and $4.0\times 10^{7}\,h^{-1}\,\mathrm{M_{\odot}}$, respectively. TNG also adopts the Planck 2015 cosmological parameters (Ade et al., 2016): $\Omega_{\Lambda,0}=0.6911$, $\Omega_{\mathrm{m},0}=0.3089$, $\Omega_{\mathrm{b},0}=0.0486$, $\sigma_{8}=0.8159$, $n_{s}=0.9667$, and $h=0.6774$.
@@ 3 The evolution of galaxy assembly bias
Here, we investigate the redshift evolution of GAB in two hydrodynamical simulations, SIMBA and TNG. Hydrodynamical simulations incorporate a wide range of astrophysical processes that might affect galaxy formation, and therefore may affect both the strength and evolution of GAB. Notably, SIMBA and TNG implement very different physical models for feedback processes, whose effects have been systematically studied by the CAMELS project (e.g., Ni et al., 2023). By studying how differences in baryonic physics lead to similarities or discrepancies in GAB evolution, we thus have the potential to identify which physical processes give rise to GAB. As a first step, our focus here is on the simpler initial question of the empirical form of GAB found in these simulations.
@@ 3.1 The total GAB
The consistent GAB trend in both simulations prior to the cosmic noon epoch ($z\sim 2$), followed by divergent behaviour at lower redshifts, is one of our key findings. This divergence coincides with the peak of the cosmic star formation rate, when feedback processes such as stellar and AGN feedback are most influential. It highlights that differences in baryonic implementations across hydrodynamical simulations can significantly affect the GAB and thus the galaxy clustering. We suspect that the difference in behaviour is tied in particular to the more explicit treatment of jets in SIMBA, which have a characteristic impact on the circumgalactic medium (Yang et al., 2024). But the detailed exploration of this issue is deferred to future work.
@@ 3.2 GAB from the cosmic environment
To investigate how much of the GAB can be attributed to the cosmic environment as described by $\delta_{5}$ or $\alpha_{5}$, we employ a double shuffling technique, during which galaxies are reassigned to haloes within the same halo mass bin as well as the $\delta_{5}$ or $\alpha_{5}$ bin. To be specific, we rank haloes in the same halo mass bin by their $\delta_{5}$ or $\alpha_{5}$ and divide the subsample into 4 quartiles accordingly. The reassigning of the galaxy contents of haloes are then performed within the same $\delta_{5}$ or $\alpha_{5}$ quartile for each halo mass bin. By doing so, the strength of GAB would decrease and the amount of change is determined by how much GAB can $\delta_{5}$ or $\alpha_{5}$ account for. For example, if all of the GAB could be attributed to $\delta_{5}$, the GAB signal after the double shuffling with $\delta_{5}$ would reduce to 0; while if none of the GAB came from $\delta_{5}$, the GAB signal after the double shuffling would remain the same level as the halo-mass only shuffling.
@@
The results for the double shuffling are shown in Fig. 3. Since the total GAB at $z=5,4,3$ is negligible in both SIMBA and TNG, we omit these redshifts from the plot. From the top panels of Fig. 3, we see that in both SIMBA and TNG, $\delta_{5}$ accounts for most of the GAB across all redshifts. The results for double shuffling with $\alpha_{5}$ are shown in the bottom panels of Fig. 3. In SIMBA, the contribution of $\alpha_{5}$ to the total GAB is negligible at all redshifts. Thus, we conclude that $\alpha_{5}$ does not contribute to GAB in SIMBA. In contrast, for TNG, $\alpha_{5}$ contributes a small but non-negligible fraction of the total GAB at all redshifts. However, this contribution remains minor compared to that from $\delta_{5}$.
@@ 4 HOD variation with cosmic environment
In the hierarchical structure formation paradigm, galaxies form within dark matter haloes, and the amount of galaxies as well as their properties are primarily governed by the halo mass. This motivates the halo occupation distribution (HOD) framework (see, e.g., Benson et al., 2000; Seljak, 2000; Peacock & Smith, 2000; White et al., 2001; Berlind & Weinberg, 2002; Cooray & Sheth, 2002), which describes the average number of galaxies hosted by haloes of a given mass. In its simplest form, the HOD assumes that halo mass alone determines galaxy occupation. In the presence of GAB, galaxy occupation is influenced not only by halo mass but also by other factors, such as the cosmic environment in which haloes reside or halo internal properties such as concentration. More recent HOD extensions introduce decorations upon the basic HOD, allowing them to incorporate GAB within the HOD framework (see e.g., Hearin et al., 2016). When there is GAB effect, we can gain valuable insights from the HOD formalism by measuring the occupation for haloes with different secondary halo properties (e.g., Artale et al., 2018; Zehavi et al., 2018). From the last section we showed that GAB can be largely attributed to the cosmic overdensity $\delta_{5}$. In this section, we further investigate how $\delta_{5}$ influences the halo–galaxy connection by examining how the HOD varies with $\delta_{5}$. For completeness, we also analyse how the HOD changes with the cosmic environment as characterised by the smoothed tidal anisotropy $\alpha_{5}$.

### v2

@@ Abstract
pubyear: 2026pagerange: Galaxy assembly bias in cosmological hydrodynamical simulations – a comparison between SIMBA and IllustrisTNGGalaxy assembly bias in cosmological hydrodynamical simulations – a comparison between SIMBA and IllustrisTNG
@@ 1 Introduction
The spatial distribution of galaxies is a key observable for cosmological studies, as it traces the underlying large-scale structure (LSS) of the Universe, which is primarily composed of invisible dark matter. In the framework of hierarchical structure formation, galaxies form within dark matter haloes that originate from primordial density fluctuations and grow through gravitational collapse, accretion, and mergers (White and Rees, 1978; Blumenthal et al., 1984; Lacey and Cole, 1993). A precise understanding of the galaxy–halo connection is therefore essential for interpreting data from current and forthcoming galaxy redshift surveys, such as the Dark Energy Spectroscopic Instrument (DESI; e.g., DESI Collaboration et al., 2024, 2025a, 2025c, 2025b) and the Euclid mission (e.g., Euclid Collaboration et al., 2022, 2025).
@@
One of the earliest approaches to modelling the galaxy–halo connection is the halo occupation distribution (HOD; e.g., Benson et al., 2000; Seljak, 2000; Peacock and Smith, 2000; White et al., 2001; Berlind and Weinberg, 2002; Cooray and Sheth, 2002), which describes the average number of galaxies hosted by a dark matter halo as a function of its mass. The simplest form of the HOD assumes that halo mass alone determines galaxy occupancy, a notion motivated by excursion set theory (Press and Schechter, 1974; Bond et al., 1991), which predicts that halo properties beyond mass are uncorrelated with the cosmic environment due to its random-walk formalism (Lemson and Kauffmann, 1999). However, this assumption was later challenged by Sheth and Tormen (2004), who demonstrated using $N$-body simulations that haloes in dense environments tend to form earlier than haloes of the same mass in less dense regions. Building on this, Gao et al. (2005) used the Millennium Simulation (Springel et al., 2005) to show that halo clustering depends not only on mass but also on halo formation time. This secondary dependence was subsequently extended to other halo properties, such as concentration and spin (Wechsler et al., 2006; Gao and White, 2007; Angulo et al., 2008; Salcedo et al., 2018; Sato-Polito et al., 2019) and the cosmic web environment (Borzyszkowski et al., 2017; Tojeiro et al., 2017; Yang et al., 2017; Paranjape et al., 2018; Han et al., 2018; Ramakrishnan et al., 2019). A dependence of halo clustering on properties beyond their mass is commonly referred to as halo assembly bias (HAB; e.g. Wechsler and Tinker, 2018), although the secondary properties may not correspond exactly to the halo assembly history.
@@
If galaxy formation is influenced by the secondary halo properties that give rise to HAB, then the core assumption of the simplest HOD model – that galaxy occupancy depends solely on halo mass – is violated. This violation can systematically alter predictions of galaxy clustering. The phenomenon in which the galaxy clustering depends on halo properties beyond mass is known as galaxy assembly bias (GAB, see Wechsler and Tinker, 2018 for a review).
@@
The existence of GAB in simulation-based catalogues was confirmed shortly after the discovery of HAB. Zhu et al. (2006) showed that galaxy properties and occupancy correlate with secondary halo properties. Croton et al. (2007) introduced the shuffling technique for quantifying GAB, and used it to demonstrate that galaxy clustering is sensitive to halo properties beyond mass. Since then, numerous studies have examined the impact of GAB on galaxy clustering and the degree to which it is driven by specific halo internal properties such as age, concentration, or spin (Zu et al., 2008; Chaves-Montero et al., 2016; Zehavi et al., 2018; Artale et al., 2018; Contreras et al., 2019; Xu et al., 2021; García-Moreno and Chaves-Montero, 2025) – or by environmental properties such as smoothed overdensity and tidal anisotropy (Zehavi et al., 2018; Artale et al., 2018; Xu et al., 2021; Alam et al., 2023; Montero-Dorta and Rodriguez, 2024; Wang et al., 2024; Perez et al., 2024; Lacerna et al., 2025). These investigations have primarily relied on galaxy catalogues generated from cosmological hydrodynamical simulations (mainly IllustrisTNG; see e.g. Montero-Dorta et al., 2020, 2021; Xu and Zheng, 2020) and/or semi-analytical galaxy formation models applied to $N$-body simulations (e.g., Contreras et al., 2019).
@@
Despite this extensive body of work, the physical nature of GAB remains unclear and its manifestation in observational data is still debated: different studies disagree over whether the driving factor is halo concentration, or on whether any GAB effect applies both to central and satellite galaxies (e.g., Vakili and Hahn, 2019; Yuan et al., 2021; Pearl et al., 2024; Zhang et al., 2025). But if GAB is indeed present in real data, then halo mass alone must be insufficient to fully describe the galaxy–halo connection in the era of precision cosmology, biasing the predicted isotropy or scale dependence of galaxy clustering (e.g. Wu et al., 2008; McCarthy et al., 2019; Obuljen et al., 2019; Hadzhiyska et al., 2020; Contreras et al., 2023). Understanding the underlying physical mechanism of GAB is therefore both critical and timely for interpreting data from galaxy redshift surveys.
@@
In this paper, we propose that comparing the strength and evolution of GAB across different cosmological hydrodynamical simulations can provide valuable insight into its physical origin. Previous studies have shown that GAB signals in IllustrisTNG and in semi-analytical models are consistent (e.g., Hadzhiyska et al., 2021), but direct comparisons between different hydrodynamical simulations have not been made. Although modern hydrodynamical simulations broadly reproduce observed galaxy properties, they adopt different subgrid prescriptions for baryonic physics – particularly stellar and AGN feedback – which can lead to substantial differences in the displacement of gas particles, halo baryon fraction, the baryonic imprint on the matter power spectrum, and star formation history (see e.g., studies of the CAMELS project; Villaescusa-Navarro et al., 2021; Ni et al., 2023; Iyer et al., 2025). Comparing GAB across simulations with differing baryonic prescriptions can thus help isolate the physical processes responsible for the emergence of GAB.
@@ 2 Data
We use two state-of-the-art cosmological hydrodynamical simulations to study the redshift evolution of GAB: SIMBA (Davé et al., 2019) and IllustrisTNG (Nelson et al., 2019). These simulations differ in their $N$-body and hydrodynamics solvers, star-formation criteria, super-massive black hole (SMBH) seeding and growth prescriptions, and most notably in their implementations of stellar and AGN feedback. These differences can lead to substantially different halo baryon fractions and intergalactic medium thermal states, which in turn can affect galaxy formation and evolution. Despite these differences, both simulations reproduce the observed galaxy stellar mass function at $z=0$ with high fidelity. Comparing the GAB evolution in the two simulations thus provides an opportunity to gain insights into the physical mechanisms driving GAB. Below, we briefly summarise the relevant features of each simulation.
@@ 2.1 SIMBA
The SIMBA simulation suite models galaxy formation within a cosmological context using a modified version of the GIZMO code (Hopkins, 2015, 2017) with the meshless finite mass (MFM) hydrodynamics solver. Stellar feedback is modelled as galactic winds. The speed of the wind particles scales with galaxy circular velocity, while the temperature of the wind particles is set by the supernovae energy minus wind kinetic energy. AGN feedback from black holes has two modes: high mass loading outflows in the radiative ‘QSO’ mode, and lower mass loading but faster outflows in the bipolar jet mode at low Eddington ratio (see Davé et al., 2019; Villaescusa-Navarro et al., 2021 and references therein).
@@ 2.2 IllustrisTNG
The IllustrisTNG project comprises a suite of large-volume cosmological magnetohydrodynamical simulations (Weinberger et al., 2017; Pillepich et al., 2018) based on the moving-mesh code AREPO (Springel, 2010). Galactic winds driven by stellar feedback are modelled kinetically as decoupled particles that are stochastically and isotropically ejected from star-forming gas, with only 10% of the supernovae energy turns into thermal energy. AGN feedback has three modes: thermal, kinetic, and radiative. Thermal mode at high accretion rate heats up the feedback sphere; kinetic mode at low accretion rate injects kinetic energy in a random direction into the feedback sphere; and the radiative mode modifies the heating and cooling of the gas surrounding the SMBH (see Weinberger et al., 2017; Pillepich et al., 2018; Villaescusa-Navarro et al., 2021 and references therein).
We use the largest run of the IllustrisTNG project, namely TNG300-1 (Nelson et al., 2019; Pillepich et al., 2017), which we refer to as TNG hereafter for simplicity. TNG simulates a cubic volume with $L_{\mathrm{box}}=205\,h^{-1}{\rm Mpc}$ using $2500^{3}$ particles for both dark matter and baryons. The mass resolution for baryon and dark matter are $7.6\times 10^{6}\,h^{-1}\,\mathrm{M_{\odot}}$ and $4.0\times 10^{7}\,h^{-1}\,\mathrm{M_{\odot}}$, respectively. TNG also adopts the Planck 2015 cosmological parameters (Ade et al., 2016): $\Omega_{\Lambda,0}=0.6911$, $\Omega_{\mathrm{m},0}=0.3089$, $\Omega_{\mathrm{b},0}=0.0486$, $\sigma_{8}=0.8159$, $n_{s}=0.9667$, and $h=0.6774$.
@@ 3 The evolution of galaxy assembly bias
Here, we investigate the redshift evolution of GAB in two hydrodynamical simulations, SIMBA and TNG. Hydrodynamical simulations incorporate a wide range of astrophysical processes that might affect galaxy formation, and therefore may affect both the strength and evolution of GAB. Notably, SIMBA and TNG implement very different physical models for feedback processes, whose effects have been systematically studied by the CAMELS project (e.g., Villaescusa-Navarro et al., 2021; Ni et al., 2023). By studying how differences in baryonic physics lead to similarities or discrepancies in GAB evolution, we thus have the potential to identify which physical processes give rise to GAB. As a first step, our focus here is on the simpler initial question of the empirical form of GAB found in these simulations.
@@ 3.1 The total GAB
The consistent GAB trend in both simulations prior to the cosmic noon epoch ($z\sim 2$), followed by divergent behaviour at lower redshifts, is one of our key findings. In both SIMBA and TNG, this redshift corresponds to the onset of substantial AGN feedback that causes galaxies to quench; prior to this redshift, feedback is dominated by stellar energy input. This suggests that AGN feedback is a key difference between these models that drives differences in the GAB and thus galaxy clustering. Indeed, AGN feedback in different models has been shown to have dramatically different impacts on the circum-galactic medium in massive haloes. We have seen that AGN feedback in SIMBA is included as stably bipolar and decoupled jets, whereas in IllustrisTNG the jet direction is randomized at every timestep and not decoupled. The result is that SIMBA produces significantly lower baryon fractions particularly in group-scale haloes (Appleby et al., 2021; Oppenheimer et al., 2021; Sorini et al., 2022), and transports gas up to $\sim 10\,h^{-1}{\rm Mpc}$ away from its origin (Borrow et al., 2020; Ni et al., 2023). This has an impact within the CGM on for instance the Sunyaev-Zeldovich $y$ decrement (Yang et al., 2022) and the intergalactic medium as traced by the Lyman alpha forest (Tillman et al., 2023). Our work here suggests that there may also be a significant impact on GAB and hence on galaxy clustering. We leave a more detailed investigation into the exact process by which AGN feedback impacts the GAB for future work.
@@ 3.2 GAB from the cosmic environment
To investigate how much of the GAB can be attributed to the cosmic environment as described by $\delta_{5}$ or $\alpha_{5}$, we employ a double-shuffling technique, during which galaxies are reassigned to haloes within the same halo mass bin as well as the $\delta_{5}$ or $\alpha_{5}$ bin. To be specific, we rank haloes in the same halo mass bin by their $\delta_{5}$ or $\alpha_{5}$ and divide the subsample into 4 quartiles accordingly. The reassigning of the galaxy contents of haloes are then performed within the same $\delta_{5}$ or $\alpha_{5}$ quartile for each halo mass bin. We now define the double-property (i.e. halo mass and $\delta_{5}$ or $\alpha_{5}$]) shuffled sample as the shuffled catalogue. This catalogue contains no GAB other than that correlated with $\delta_{5}$ or $\alpha_{5}$. The additional GAB signal in the original sample, quantified by $w_{\mathrm{p}}(r_{\mathrm{p}})/w_{\mathrm{p}}(r_{\mathrm{p}})_{\mathrm{shuffled}}-1]$, would therefore reduce to zero if the double-property shuffled sample retained the same level of GAB as the original sample. Conversely, if no GAB were associated with $\delta_{5}$ or $\alpha_{5}$, the double-property shuffled sample would contain no GAB, and the resulting $w_{\mathrm{p}}(r_{\mathrm{p}})/w_{\mathrm{p}}(r_{\mathrm{p}})_{\mathrm{shuffled}}-1]$ would be identical to that obtained from the halo-mass-only shuffled sample.
@@
The results for the double shuffling are shown in Fig. 3. Since the total GAB at $z=5,4,3$ is negligible in both SIMBA and TNG, we omit these redshifts from the plot. From the top panels of Fig. 3, we see that in both SIMBA and TNG, $\delta_{5}$ accounts for most of the GAB across all redshifts. The results for double shuffling with $\alpha_{5}$ are shown in the bottom panels of Fig. 3. In SIMBA, the contribution of $\alpha_{5}$ to the total GAB is negligible at all redshifts. Thus, we conclude that $\alpha_{5}$ does not contribute to GAB in SIMBA. In contrast, for TNG, $\alpha_{5}$ contributes a small but non-negligible fraction of the total GAB at all redshifts. However, this contribution remains minor compared to that from $\delta_{5}$. Notice that a dependence of GAB on $\delta_{5}$ implies that halo assembly bias (HAB) must itself depend on $\delta_{5}$, as demonstrated in previous studies (e.g. Dalal et al., 2008; Tucci et al., 2020; Mansfield and Kravtsov, 2020). When this HAB dependence is combined with a tendency for galaxy formation to vary with $\delta_{5}$, the result is a GAB dependence on $\delta_{5}$. We would further explore this tendency in the next Section.
@@ 4 HOD variation with cosmic environment
In the hierarchical structure formation paradigm, galaxies form within dark matter haloes, and the amount of galaxies as well as their properties are primarily governed by the halo mass. This motivates the halo occupation distribution (HOD) framework (see, e.g., Benson et al., 2000; Seljak, 2000; Peacock and Smith, 2000; White et al., 2001; Berlind and Weinberg, 2002; Cooray and Sheth, 2002), which describes the average number of galaxies hosted by haloes of a given mass. In its simplest form, the HOD assumes that halo mass alone determines galaxy occupation. In the presence of GAB, galaxy occupation is influenced not only by halo mass but also by other factors, such as the cosmic environment in which haloes reside or halo internal properties such as concentration. More recent HOD extensions introduce decorations upon the basic HOD, allowing them to incorporate GAB within the HOD framework (see e.g., Hearin et al., 2016). When there is GAB effect, we can gain valuable insights from the HOD formalism by measuring the occupation for haloes with different secondary halo properties (e.g., Artale et al., 2018; Zehavi et al., 2018). From the last section we showed that GAB can be largely attributed to the cosmic overdensity $\delta_{5}$. In this section, we further investigate how $\delta_{5}$ influences the halo–galaxy connection by examining how the HOD varies with $\delta_{5}$. For completeness, we also analyse how the HOD changes with the cosmic environment as characterised by the smoothed tidal anisotropy $\alpha_{5}$.